OpenStudy (geometry_hater):

Side one of a triangle is (3x+10), side two of a triangle is (5x) and side 3 of a triangle is (2x + 25). If the perimeter of the triangle is 200, find the value of x. Side one of a triangle is (3x+10), side two of a triangle is (5x) and side 3 of a triangle is (2x + 25). If the perimeter of the triangle is 200, find the value of x. @Mathematics

hero (hero):

3x+10 + 5x + 2x+25 = 200 10x + 35 = 200 10x = 200 - 35 10x = 165 x = 16.5